# A Hybrid Quantum Neural Network for Split Learning

### v1

@@ 6.1 Reconstruction Attack from Smashed Data in HQSL
In this section, we present our threat model that describes how an honest but curious server adversary can recreate the private input raw data from the smashed data transmitted by the client-side model. The adversary uses a reconstruction attack model to recreate the client’s input data. Hence, we subsequently describe three such reconstruction attack models that can be employed by the adversary. We then describe how we design our proposed defense mechanism to mitigate the risk of reconstruction attacks in HQSL. This countermeasure entails the introduction of a Laplacian noise layer at the end of the client-side model, designed and configured considering the rotational properties of encoding gates in the server-side quantum layer.
@@ 6.1.3 Encoding Gate-based Noise Layer Defense Mechanism
To strengthen the security of HQSL against the risk of reconstruction attacks, we propose a Laplacian noise layer defense mechanism at the end of the client-side model. Previous works have considered the use of a noise layer defense mechanism to introduce perturbation to the smashed data and hence, increase the difficulty of reconstruction by the proposed reconstruction models. However, such noise defense methods come with a privacy-utility trade-off [33, 20, 17]. In this paper, we study the rotational properties of quantum encoding gates to design the noise layer to (i) effectively hinder the reconstruction attack, and (ii) maintain a high classification performance by the server during inference. This, hence, addresses the privacy-utility trade-off associated with the noise layer defense mechanism. We empirically compare these two aspects for the hybrid (HQSL) and classical (SL) cases in Sections 6.2 and 6.3.
Laplacian noise is favored over other types of noise, such as Gaussian noise, for privacy protection or data obfuscation. This is because noise sampled from a Laplacian distribution, $Laplace(\mu,b)$, can be systematically added to the smashed data, $\textsf{Z}_{\textsf{priv}}$, to provide controlled randomness, while enabling the calculation of the level of perturbation introduced. This is important for maintaining privacy while retaining data utility. Laplacian noise is also commonly utilized in the context of differential privacy [29, 28].The location or mean parameter, $\mu$, provides a translational property to the data, while the scale parameter, $b$, indicates the variance or perturbation level caused by the noise. The scale parameter also determines the sharpness of the Laplacian distribution about its mean value as shown in Fig. 10(a). The smaller the scale parameter, the sharper the Laplacian distribution is about its mean value, hence sampling values close to the mean.
@@
(b) Rotation of qubit (red dot) around the x-axis of the Bloch sphere after adding Laplacian noise, $\eta$, with mean $\mu=4\pi$ and small scale, $b$
Figure 10: (a) As the scale value $b$ decreases, the Laplacian noise becomes more concentrated around the mean, $\mu$, resulting in a higher density of data points near $\mu$. (b) The addition of this noise (mean $\mu=4\pi$ and small scale $b$) to the smashed data causes only a small fluctuation near the original location (red dot) of the qubit about the x-axis on the Bloch sphere
@@
Adding noise sampled from a Laplacian distribution, with appropriately tuned mean and scale parameters, has the potential to benefit the quantum circuit on the server side in terms of maintaining a high classification performance and impeding reconstruction of raw data due to the jitter added to the data. In Section 6.3, we will demonstrate how these are not possible for a server without a quantum layer. We consider the following to develop a better understanding of our selection of Laplacian noise parameters.
@@
The Bloch sphere gives a geometrical representation of the pure state space of a qubit. Quantum gates, e.g., single-qubit rotation gates cause a qubit’s state vector to move around the surface of the Bloch sphere as shown in Fig. 10(b). In our proposed quantum circuit for HQSL (Fig. 2(b)), RX-gates have a period of rotation of $4\pi$ and are responsible for encoding classical data to quantum states. These single-qubit rotation gates rotate the state vector of qubit around the X-axis of the Bloch sphere as per an angle parameter, $\theta$, equal to the numerical value of each classical input feature, $z$, in our HQSL model. We can thus write $RX(\theta)=RX(z)$. An additive noise to the classical smashed data corresponds to a noise, $\eta$, that causes rotation of the qubit by a perturbed angle, $\tilde{\theta}=z+\eta$, about the x-axis of the Bloch sphere. Hence, $RX(\tilde{\theta})=RX(z+\eta)$. Furthermore, if $\eta$ follows a Laplacian distribution with $\mu=4k\pi$, for $k\in\mathbb{Z}$ and a small scale $b$, the resulting quantum state is similar to that before the noise is applied. This is due to the rotational property of our encoding gates used in the quantum layer of HQSL. We formalize this below.
In the hybrid setting, for $\mu=4k\pi$, $k\in\mathbb{Z}$, i.e., $\mu\in\{\dots,-4\pi,0,4\pi,8\pi,\dots\}$, we have $\eta$ $\sim$ $Laplace(\mu,b)$. Equivalently, due to the periodicity of the RX-gate, $\eta$ $\sim$ $Laplace(0,b)$, and we have
$\lim_{b\to 0}RX(z+\eta)=RX(z).$ (7)
Next, if we consider a mean of $\mu=(4k+2)\pi$, where $k\in\mathbb{Z}$, i.e., $\mu\in\{\dots,-6\pi,-2\pi,2\pi,6\pi,\dots\}$, the resulting quantum state encoded by the RX-gate corresponds to a global phase factor of $-1$ to the noise-free quantum state. Hence, for $\mu=(4k+2)\pi$, we can update Eq. 7 as:
$\lim_{b\to 0}RX(z+\eta)=-RX(z).$ (8)
In quantum mechanics, a global phase factor of -1 does not affect the physical state. This means that $RX(z+\eta)$, where $\eta\sim Laplace(\mu,b)$ and $\mu=(4k+2)\pi$, is equivalent to $RX(z)$ up to a global phase. This does not affect the further processing along the circuit or measurement outcomes, as it remains a global factor throughout the entire circuit evolution [19].
Hence, in general, for the quantum layer on the server side, given $\mu=2k\pi$, where $k\in\mathbb{Z}$, and a small enough scale parameter $b$, the quantum state encoded in the quantum circuit from the noisy smashed data, $\tilde{\textsf{Z}}_{\textsf{priv}}$, is expected to be close to the quantum state corresponding to its clean version, $\textsf{Z}_{\textsf{priv}}$. This ensures that the performance of the server model in HQSL is maintained. In contrast, due to the classical nature of the reconstruction model and the server model in SL, the additive noise can adversely impact their performances. We demonstrate these in Section 6.3.
@@
Moreover, maintaining a small scale parameter, $b$, allows us to limit the noise to a minimal perturbation about the mean, $\mu$, as indicated by the sharpness of the Laplacian PDF at its mean value (Fig. 10(a)). This ensures a low variability in the resulting quantum state, allowing a consistent encoding that closely corresponds to the clean version of the quantum state. Thus, the primary objective for keeping the scale parameter, $b$, small is to preserve the utility of the hybrid model.
In Fig. 11, we illustrate the defense setup deployed at inference time after training the HQSL/SL models. $\textsf{X}_{\textsf{inf}}$ and $\textsf{Z}_{\textsf{inf}}$ represent the client’s private input data and smashed data from the client model at inference time, respectively. The defense mechanism involves using the Laplacian noise layer to obscure the clean smashed data, $\textsf{Z}_{\textsf{inf}}$, thereby making it more challenging for the attack models to reconstruct the input data. $\tilde{\textsf{Z}}_{\textsf{inf}}$ represents the noisy version of $\textsf{Z}_{\textsf{inf}}$.
@@ 6.2 Experiments on Encoding Gate-Based Noise Defense
We use the stochastic gradient descent, SGD, optimizer, with a learning rate of $10^{-3}$ and momentum of $0.9$, to train the reconstruction model for 200 epochs using the MAE/L1Loss loss function. After training the adversary’s reconstruction model, we introduce the Laplacian noise layer, with mean $\mu\in\{0,\pi,2\pi,3\pi,4\pi\}$ and scale $b\in\{0.01,0.1,0.5,1\}$, at the end of the previously trained client-side model. We consider this range of values for the mean $\mu$ because they represent the location of a qubit at different points (including the cases of $\mu=2\pi$ and $4\pi$ expected to cause the quantum states to be invariant in the presence of our noise layer given a small scale $b$) in a full rotation around the x-axis of the Bloch sphere.
$\textsf{X}_{\textsf{inf}}$ represents a portion of the private data of a client at inference time, $\textsf{X}_{\textsf{inf}}\subset\mathcal{D}_{\textsf{priv}}$, that we aim to reconstruct using the trained reconstruction model. However, due to the noise layer, the client model now outputs noisy smashed data, $\tilde{\textsf{Z}}_{\textsf{inf}}$. $\tilde{\textsf{Z}}_{\textsf{inf}}$ represents the input to the reconstruction model during the attack phase to recreate $\textsf{X}_{\textsf{inf}}$ as $\hat{\mathbf{\textsf{X}}}_{\textsf{inf}}$.
@@
For evaluating the classification performance at inference time, we compare the test accuracies and F1-scores under HQSL and SL settings for five-folds, now in the presence of the Laplacian noise layer with parameters $\mu$ and $b$. These results are discussed in Section 6.3.2.
@@ 6.3 Reconstruction Attacks Results and Discussion
In this section, we present and discuss our findings from the experiments carried out to test our proposed encoding gate-based noise defense layer as a countermeasure against reconstruction attacks. We evaluate the performance of the reconstruction model by using the metrics described earlier in 6.3.1. This is done in both classical and hybrid settings and for different Laplacian noise parameter values, $\mu$ and $b$. Next, we present the inference time test accuracies and F1-scores of HQSL against SL across five folds by fixing the mean parameter, $\mu$, and varying the scale parameter, $b$, in 6.3.2. These experimental findings allow us to tune the noise parameters to give HQSL an advantage over SL in terms of protection against reconstruction attacks and inference time classification performance despite the presence of the noise layer.
@@
To evaluate the performance of the reconstruction model in classical and hybrid settings in the presence of Laplacian noise, we investigated the impact of noise mean, $\mu$, and scale, $b$, in recreating Fashion-MNIST, MNIST, and spectrogram images. Our comparative analysis focused on 4 key image comparison metrics: Mean Cosine Distance, Mean MSE, Mean Structural Dissimilarity Index (Mean DSSIM), and Mean Log Spectral Distance (Mean LSD). The results for the FMNIST, MNIST, and Speech Commands spectrograms datasets are depicted in Fig. 12(a), 12(a) and 12(a) respectively. The baseline values (represented by the horizontal dashed lines) correspond to the noise-free implementation of the reconstruction attack, i.e., when the smashed data transmitted to the reconstruction model are free from Laplacian noise. In both the classical and hybrid settings, the baseline results are closer to 0 than when the noise layer is applied, showing the reconstruction attacks are successful in recreating the private inputs. These results, hence, underscore the need for the defense mechanism that we investigate in this section.
@@
The general trends from Fig. 12(a) and 12(a) were that as we increased the mean parameter, $\mu$, the metric values were well above the baseline, irrespective of the noise scale parameter, $b$. Notably, on the MNIST dataset, the deviations from the baseline in the hybrid setting were more significant than in the classical setting for all 3 metrics. This shows that there was a larger difference between original and reconstructed images in the hybrid setting than in the classical setting. In the classical setting, the minimal deviations from the baseline indicate that a reconstruction attack can still be successful for the MNIST dataset, regardless of the noise parameters. This could reflect a model characteristic where the noise parameters do not perturb the foundational performance of the reconstruction model. We also highlight that at $\mu=4\pi$, irrespective of the scale parameter value, $b$, the largest deviations were obtained, particularly in the hybrid setting, indicating that at this $\mu$ value, the additive noise was successfully hindering the reconstruction performance of the adversary’s model. We also note that significant deviations were still obtained at $\mu=2\pi$. These findings underscore that the mean values $\mu=2\pi$ and $4\pi$ caused the performance of the adversary’s reconstruction model to drop in the hybrid setting.
@@
Overall, this study demonstrates that the introduction of the Laplacian noise layer, with appropriately configured noise parameters, can obfuscate the smashed data thereby hindering the performance of the reconstruction model, and in some cases, to a much greater extent in the hybrid setting. From these results, we establish that mean values of $\mu=2\pi$ and $4\pi$ effectively increase the difference between the original and reconstructed images, hence supporting our analysis in Section 6.1.3. Therefore, we conclude here that by tuning the Laplacian noise layer parameters considering the rotational properties of encoding gates, we can strengthen the security of HQSL against reconstruction attacks. In the classical setup, the reconstruction attack can still successfully recreate private input data. Next, a necessary investigation is the impact of the perturbation caused by the scale parameter, $b$ on the classification performance of HQSL and SL.

### v2

@@ 6.1 Reconstruction Attack from Smashed Data in HQSL
In this section, we present our threat model that describes how an honest but curious server adversary can recreate the private input raw data from the smashed data transmitted by the client-side model. The adversary uses a reconstruction attack model to recreate the client’s input data at inference time. Hence, we subsequently describe three such reconstruction attack models that can be employed by the adversary. We then describe how we design our proposed defense mechanism to mitigate the risk of reconstruction attacks in HQSL. This countermeasure entails the introduction of a Laplacian noise layer at the end of the client-side model, designed and configured considering the rotational properties of encoding gates in the server-side quantum layer.
@@ 6.1.3 Encoding Gate-based Noise Layer Defense Mechanism
To defend HQSL against the risk of reconstruction attacks, we propose a Laplacian noise layer defense mechanism at the end of the client-side model. Previous works have considered the use of a noise layer defense mechanism to introduce perturbation to the smashed data and hence, increase the difficulty of reconstruction by the proposed reconstruction models. However, such noise defense methods come with a privacy-utility trade-off [37, 24, 21]. In this section, we study the rotational properties of quantum encoding gates to design the noise layer to (i) effectively hinder the reconstruction attack, and (ii) maintain a high classification performance (accuracy and F1-score) by the server during inference. This, hence, addresses the privacy-utility trade-off associated with the noise layer defense mechanism. We empirically compare these two aspects for the hybrid (HQSL) and classical (SL) cases in Sections 6.2 and 6.3.
Laplacian noise is favored over other types of noise, such as Gaussian noise, for privacy protection or data obfuscation. This is because noise sampled from a Laplacian distribution, $\mathrm{Laplace}(\mu,b)$, can be systematically added to the smashed data, $\textsf{Z}_{\textsf{priv}}$, to provide controlled randomness, while enabling the explicit quantification of the level of perturbation introduced. This is important for maintaining privacy while retaining data utility. Laplacian noise is also commonly utilized in the context of differential privacy [33, 32].
The location (mean) parameter $\mu$ shifts the distribution, while the scale parameter $b$ controls its spread. In our setting, this is critical: the scale parameter determines how tightly noise samples concentrate around the mean. As shown in Fig. 11(a), smaller $b$ values result in a sharper distribution centered at $\mu$, keeping the noise values sampled close to the desired noise level. This sharpness allows us to maintain high fidelity in the quantum state encoding while still introducing enough variability to hinder adversarial reconstruction, as we show next.
@@
(b) Rotation of qubit (red dot) around the x-axis of the Bloch sphere after adding Laplacian noise, $\eta$, with mean $\mu=4k\pi$ and small scale, $b$
Figure 11: (a) As the scale value $b$ decreases, the Laplacian noise becomes more concentrated around the mean, $\mu$, resulting in a higher density of data points near $\mu$. (b) The addition of this noise (mean $\mu=4k\pi$, $k\in\mathbb{Z}$, and small scale $b$) to the smashed data causes only a small fluctuation near the original location (red dot) of the qubit about the x-axis on the Bloch sphere
@@
Adding noise sampled from a Laplacian distribution, with appropriately tuned mean and scale parameters, has the potential to benefit the quantum circuit on the server side by maintaining high classification performance while impeding reconstruction of raw data via the added jitter. In Section 6.3, we will demonstrate that a purely classical server cannot simultaneously achieve these goals.
@@
We consider the following to develop a better understanding of our selection of Laplacian noise parameters. The Bloch sphere gives a geometrical representation of the pure‐state space of a qubit. Single‐qubit rotation gates, e.g., the RX-gate, move the state vector around the surface of the Bloch sphere, as shown in Fig. 11(b). In our proposed HQSL circuit (Fig. 2(b)), each RX‐gate has a $4\pi$ period and encodes a classical feature $z$ by rotating the qubit about the x‐axis by angle $\theta=z$. Adding noise $\eta$ to the smashed data, hence, $\tilde{\theta}=z+\eta$, corresponds to applying
$RX(\tilde{\theta})\;=\;RX(z+\eta)\,.$ (7)
By the $4\pi$ periodicity of $RX(\theta)$, centering our Laplace noise at any mean $\mu=2k\pi$ ($k\in\mathbb{Z}$) ensures that as $b\to 0$,
$RX(z+\eta)\;\approx\;RX(z)\quad(\text{up to a global phase}),$ (8)
so that the encoded quantum state is preserved. By computing the fidelity between the clean and noisy encoded quantum states, we formalize this below.
Expected Fidelity Preservation.
Using the matrix representation of our encoding RX-gate, the encoded quantum state, $\lvert\psi(z)\rangle$, can be written as follows:
$\lvert\psi(z)\rangle=RX(z)\lvert 0\rangle=\cos\!\tfrac{z}{2}\lvert 0\rangle-i\sin\!\tfrac{z}{2}\lvert 1\rangle,$ (9)
and with noise $\eta\sim\mathrm{Laplace}(0,b)$ we have:
$\lvert\psi(z+\eta)\rangle=\cos\!\tfrac{z+\eta}{2}\lvert 0\rangle-i\sin\!\tfrac{z+\eta}{2}\lvert 1\rangle.$ (10)
The state fidelity is
$F(\eta)=\bigl{|}\langle\psi(z)\mid\psi(z+\eta)\rangle\bigr{|}^{2}=\cos^{2}\!\tfrac{\eta}{2}=\tfrac{1+\cos\eta}{2}.$ (11)
Taking expectation over $\eta$ and using the Laplace characteristic function yields
$\mathbb{E}_{\eta}[F]=\frac{1}{2}\Bigl{(}1+\mathbb{E}[\cos\eta]\Bigr{)}=\frac{1}{2}\Bigl{(}1+\tfrac{1}{1+b^{2}}\Bigr{)}.$ (12)
@@
In particular, for $b\ll 1$, we have $\mathbb{E}_{\eta}[F]\approx 1$. This implies that the clean and noisy encoded quantum state on the server-side quantum layer remains unchanged as $b\to 0$. Hence, keeping the scale parameter $b$ small limits the noise to a minimal perturbation about the mean, as illustrated by the sharp Laplacian PDF in Fig. 11(a). This low variability ensures consistent encoding that closely matches the clean quantum state, preserving the hybrid server model’s utility.
In Fig. 12, we illustrate the defense setup at inference time: the client’s smashed data $\textsf{Z}_{\textsf{inf}}$ is perturbed to $\tilde{\textsf{Z}}_{\textsf{inf}}$ via the Laplacian noise layer before entering the server’s quantum circuit, thereby obscuring the input against reconstruction attacks while retaining classification accuracy.
@@
We also include Algorithm 3 to outline the deployment of the Laplacian noise-based defense mechanism at inference time as a countermeasure against reconstruction attacks on the server side. The algorithm describes the process of obfuscating the smashed data $\textsf{Z}_{\texttt{inf}}$ before sending $\tilde{\textsf{Z}}_{\texttt{inf}}$ to the HQSL server. Depending on the selected mode, the server either performs classification using the hybrid quantum server model or attempts to reconstruct the original input $\textsf{X}_{\texttt{inf}}$.
Input: Pre-trained Client Model $f_{c}$ with weights $\mathbf{\tilde{W}}^{C}$, Server Model with classical layers $f_{s}$ and weights $\mathbf{\tilde{W}}^{S}$, quantum layer parameters $\mathbf{\tilde{\Theta}}$, Adversary Reconstruction Model $f_{\texttt{rec}}$, Input Data $\textsf{X}_{\texttt{inf}}$, Mode Selection $mode$ $(classification/attack)$
Output: Classification Output $\hat{\mathbf{\textsf{Y}}}_{\texttt{inf}}$ (if classification mode) or Adversary’s Reconstructed Input $\hat{\mathbf{\textsf{X}}}_{\texttt{inf}}$ (if attack mode)
Notations: Client-side smashed data: $\textsf{Z}_{\texttt{inf}}$, quantum layer expectation output: $\mathbf{E}_{\texttt{inf}}$
START OF INFERENCE
Client Side: // Forward Pass
Compute smashed data: $\textsf{Z}_{\texttt{inf}}=f_{c}(\textsf{X}_{\texttt{inf}},\tilde{\mathbf{W}}^{C})$
Sample Laplacian noise: $\eta\sim\text{Laplace}(\mu,b)$ // Noise sampled from Laplace distribution with mean $\mu$ and scale $b$
Add noise to smashed data: $\tilde{\textsf{Z}}_{\texttt{inf}}\leftarrow\textsf{Z}_{\texttt{inf}}+\eta$ // Obfuscate smashed data
Send $\tilde{\textsf{Z}}_{\texttt{inf}}$ to the server
Server Side: // Either Classification or Adversary Attack
if $mode=classification$ then
Compute expectation values: $\mathbf{E}_{\texttt{inf}}=f_{q}(\tilde{\textsf{Z}}_{\texttt{inf}},\tilde{\mathbf{\Theta}})$ // Quantum Layer
Compute classification output: $\hat{\mathbf{\textsf{Y}}}_{\texttt{inf}}=f_{s}(\mathbf{E}_{\texttt{inf}},\tilde{\mathbf{W}}^{S})$ // Classical Layers
return $\hat{\mathbf{\textsf{Y}}}_{\texttt{inf}}$
else if $mode=attack$ then
Attempt to reconstruct input: $\hat{\mathbf{\textsf{X}}}_{\texttt{inf}}\leftarrow f_{\texttt{rec}}(\tilde{\textsf{Z}}_{\texttt{inf}})$ // Adversary Reconstruction
return $\hat{\mathbf{\textsf{X}}}_{\texttt{inf}}$
Main Execution:
Load input data $\textsf{X}_{\texttt{inf}}$ and compute smashed data $\textsf{Z}_{\texttt{inf}}$
Compute $\tilde{\textsf{Z}}_{inf}\leftarrow$ Sample and add Laplacian noise $\eta$ to $\textsf{Z}_{\texttt{inf}}$
if $mode=classification$ then $\hat{\mathbf{\textsf{Y}}}_{\texttt{inf}}\leftarrow$ Inference($\tilde{\textsf{Z}}_{\texttt{inf}}$)
else if $mode=attack$ then $\hat{\mathbf{\textsf{X}}}_{\texttt{inf}}\leftarrow$ Adversary Reconstruction($\tilde{\textsf{Z}}_{\texttt{inf}}$)
END OF INFERENCE
return $\hat{\mathbf{\textsf{Y}}}_{\texttt{inf}}$ (if classification mode) or $\hat{\mathbf{\textsf{X}}}_{\texttt{inf}}$ (if attack mode)
Algorithm 3 HQSL with Laplacian Noise Defense Deployed at Inference Time
@@ 6.2 Experiments on Encoding Gate-Based Noise Defense
We use the stochastic gradient descent, SGD, optimizer, with a learning rate of $10^{-3}$ and momentum of $0.9$, to train the reconstruction model for 200 epochs using the MAE/L1Loss loss function. After training the adversary’s reconstruction model, we introduce the Laplacian noise layer, with mean $\mu\in\{0,\pi,2\pi,3\pi,4\pi\}$ and scale $b\in\{0.01,0.1,0.5,1\}$, at the end of the previously trained client-side model. We consider this range of values for the mean $\mu$ because it represents the location of a qubit at different points along a full rotation around the x-axis of the Bloch sphere. Notably, due to the $4\pi$-periodicity of quantum states under such rotations, angles beyond $4\pi$ do not provide additional insights while tuning our Laplacian noise defence layer, as they result in equivalent quantum states. For example, mean values $\mu=6\pi$ and $8\pi$ would be redundant as they are equivalent to $\mu=2\pi$ and $4\pi$ respectively, as the state evolution repeats cyclically with a period of $4\pi$. The cases of $\mu=2\pi$ and $4\pi$ are particular relevant, as they are expected to cause the quantum states to be invariant in the presence of our noise layer given a small scale $b$ as explained in Section 6.1.3.
Next, $\textsf{X}_{\textsf{inf}}$ represents a portion of the private data of a client at inference time, $\textsf{X}_{\textsf{inf}}\subset\mathcal{D}_{\textsf{priv}}$, that we aim to reconstruct using the trained reconstruction model. However, due to the noise layer, the client model now outputs noisy smashed data, $\tilde{\textsf{Z}}_{\textsf{inf}}$. $\tilde{\textsf{Z}}_{\textsf{inf}}$ represents the input to the reconstruction model during the attack phase to recreate $\textsf{X}_{\textsf{inf}}$ as $\hat{\mathbf{\textsf{X}}}_{\textsf{inf}}$.
@@
For evaluating the classification performance at inference time, we compare the test accuracies and F1-scores under HQSL and SL settings for five folds, now in the presence of the Laplacian noise layer with parameters $\mu$ and $b$. These results are discussed in Section 6.3.2.
@@ 6.3 Reconstruction Attacks Results and Discussion
In this section, we present and discuss our findings from the experiments carried out to test our proposed encoding gate-based noise defense layer as a countermeasure against reconstruction attacks. We evaluate the performance of the reconstruction model by using the metrics described earlier in 6.3.1. This is done in both classical and hybrid settings and for different Laplacian noise parameter values, $\mu$ and $b$. Next, we analyze HQSL’s inference-time accuracies and F1-scores across five folds, fixing the mean parameter $\mu$ while varying the scale parameter $b$, as detailed in Section 6.3.2. These findings provide insights into optimizing the noise parameters to enhance HQSL’s defense against reconstruction attacks while maintaining robust inference-time classification performance compared to SL, despite the added noise layer.
@@
To evaluate the performance of the reconstruction model in classical and hybrid settings in the presence of Laplacian noise, we investigated the impact of noise mean, $\mu$, and scale, $b$, in reconstructing Fashion-MNIST, MNIST, and spectrogram images. Our comparative analysis focused on 4 key image comparison metrics: Mean Cosine Distance, Mean MSE, Mean Structural Dissimilarity Index (Mean DSSIM), and Mean Log Spectral Distance (Mean LSD). The results for the FMNIST, MNIST, and Speech Commands spectrograms datasets are depicted in Fig. 13(a), 13(a) and 13(a) respectively. The baseline values (represented by the horizontal dashed lines) correspond to the noise-free implementation of the reconstruction attack, i.e., when the smashed data transmitted to the reconstruction model are free from Laplacian noise. In both the classical and hybrid settings, the baseline results are closer to 0 than when the noise layer is applied, showing the reconstruction attacks are successful in recreating the private inputs. These results, hence, underscore the need for the defense mechanism that we investigate in this section.
@@
The general trends from Fig. 13(a) and 13(a) were that as we increased the mean parameter, $\mu$, the metric values were well above the baseline, irrespective of the noise scale parameter, $b$. Notably, on the MNIST dataset, the deviations from the baseline in the hybrid setting were more significant than in the classical setting for all 3 metrics. This shows that there was a larger difference between original and reconstructed images in the hybrid setting than in the classical setting. In the classical setting, the minimal deviations from the baseline indicate that a reconstruction attack can still be successful for the MNIST dataset, regardless of the noise parameters. This suggests that, at least for the MNIST dataset, the reconstruction model in the classical setting appears relatively insensitive to changes in the noise parameters, with its baseline performance being largely preserved regardless of the trialled noise configurations. We also highlight that at $\mu=4\pi$, irrespective of the scale parameter value, $b$, the largest deviations were obtained, particularly in the hybrid setting, indicating that at this $\mu$ value, the additive noise was successfully hindering the reconstruction performance of the adversary’s model. We also note that significant deviations were still obtained at $\mu=2\pi$. These findings underscore that the mean values $\mu=2\pi$ and $4\pi$ caused the performance of the adversary’s reconstruction model to drop in the hybrid setting.
@@
Overall, this study demonstrates that the introduction of the Laplacian noise layer, with appropriately configured noise parameters, can obfuscate the smashed data thereby hindering the performance of the reconstruction model, and in some cases, to a much greater extent in the hybrid setting. From these results, we establish that mean values of $\mu=2\pi$ and $4\pi$ effectively increase the difference between the original and reconstructed images, hence supporting our analysis in Section 6.1.3. Therefore, we conclude here that by tuning the Laplacian noise layer parameters considering the rotational properties of encoding gates, we can strengthen the resilience of HQSL against reconstruction attacks. In the classical setup, the reconstruction attack can still successfully recreate private input data. Next, a necessary investigation is the impact of the perturbation caused by the scale parameter, $b$ on the classification performance of HQSL and SL.